A and B have a certain number of mangoes . A says to B " If you give 30 of your mangoes I will have twice as many left with you ." B replies , " If you give me 10 , I will have thrice as many as left with you . How many mangoes does each have ? ( Pls help me )
Answer by mukhopadhyay(490)

Assume A has x mangoes and B has y mangoes.
(x+30)=2(y-30)........ (Eq 1)
=>x=2y-90
3(x-10)=(y+10)........ (Eq 2)
=>3x=y+40
=>3(2y-90) = y+40 (using x from Eq 1)
=>6y-y=310
=>y=62
So, x=124-90=34
Answer: A has 34 mangoes and B has 62 mangoes